OpenStudy (anonymous):

write the equation of a line (in slope-intercept form) that pases through (3,-4) and is parallel to y=-3x 2

OpenStudy (anonymous):

y-y1 = m(x-x1) m = -3, according to the given info (3, -4) (x1, y1) plug the points

OpenStudy (anonymous):

then put it in to "y=mx+b" form